OpenStudy (anonymous):

A light plane must reach a speed of 33 m/s per take off. How long a runway is needed if the constant acceleration is 3.0 m/s^2 Answer: 1.8x10^2 meters.

OpenStudy (anonymous):

v2^2 = v1^2 + 2ad (33)^2 = 0m/s + 2(3m/s^2) d 1089 = 6m/s^2 d d= 1089/6 181.5 m OR 1.8 x 10^2m